OpenStudy (anonymous):

rewrite the following problem in ditributive property. 5(2x-8) A)-30x B) 10x-40 C)10x+40 D)10x-8 @coolshava

OpenStudy (anonymous):

im thinking its either B or c

OpenStudy (anonymous):

yeah. the difference is that it's 5(2x) MINUS 5(8). that trips me up a lot.

OpenStudy (anonymous):

why C or B

OpenStudy (anonymous):

look to the sign

OpenStudy (anonymous):

maybe 5/2 ?

OpenStudy (anonymous):

10x-8

OpenStudy (anonymous):

whenever you're using distributive, it's usually going to be multiplying. you were right with the B or C part, you just need to look at the signs on the original problem too.

OpenStudy (anonymous):

ok 10x-40 becuase its subration?

OpenStudy (anonymous):

5(2x-8)=5*2x-5*8

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

awesome i get it now thanks!

OpenStudy (anonymous):

simplify the following expression 3(5-12)+4/4 A)19 B)-19 C)23 D)8.5 @coolshava @AJ01

OpenStudy (anonymous):

just do the same thing

OpenStudy (anonymous):

3(5-12)+4/4 3*5-3*12+4/4 15-36+1

OpenStudy (anonymous):

multiply 3 by the others?

OpenStudy (anonymous):

-19?

OpenStudy (anonymous):

15-36= -21 +1 = -20 @AJ01

OpenStudy (anonymous):

15-36+1=16-36=-20

OpenStudy (anonymous):

but that isnt one of the options? @AJ01

OpenStudy (anonymous):

i know that... if you didn't miss anything in the question then this is the answer

OpenStudy (anonymous):

\[3(5-12)+\frac{ 4 }{ 4 }=-20\]

OpenStudy (anonymous):

its not 4 4ths its 4 divided by 4 @AJ01

OpenStudy (anonymous):

rewrite it using Equation form

OpenStudy (anonymous):

i have no idea i seriously suck at math @AJ01

OpenStudy (anonymous):

ok it is like this : \[3(5-12)+4/4\] or \[\frac{ 3(5-12)+4 }{ 4}\]

OpenStudy (anonymous):

so -19? @AJ01

OpenStudy (anonymous):

the closest is -19 but still not right

OpenStudy (anonymous):

what about 23?

OpenStudy (anonymous):

non of the 2 form that i wrote would give -19

OpenStudy (anonymous):

maybe is a miss take in the book

OpenStudy (anonymous):

if the problem 3(5-12)+(4/4) then the answer is -20.